# VALLR-Pin: Dual-Decoding Visual Speech Recognition for Mandarin with Pinyin-Guided LLM Refinement

### v1

@@ Abstract
Visual Speech Recognition (VSR) aims to transcribe spoken words from silent lip-motion videos. This task is particularly challenging for Mandarin, as visemes are highly ambiguous and homophones are prevalent. We propose VALLR-Pin, a novel two-stage framework that extends the recent VALLR architecture from English to Mandarin. First, a shared video encoder feeds into dual decoders, which jointly predict both Chinese character sequences and their standard Pinyin romanization. The multi-task learning of character and phonetic outputs fosters robust visual–semantic representations. During inference, the text decoder generates multiple candidate transcripts (N-best list). We construct a prompt by concatenating the Pinyin output with these candidate Chinese sequences and feed it to a large language model (LLM) to resolve ambiguities and refine the transcription. This provides the LLM with explicit phonetic context to correct homophone-induced errors. Finally, we fine-tune the LLM on synthetic ”noisy” examples: we generate imperfect Pinyin–text pairs from intermediate VALLR-Pin checkpoints using the training data, creating instruction–response pairs for error correction. This endows the LLM with awareness of our model’s specific error patterns. In summary, VALLR-Pin synergizes visual features with phonetic and linguistic context to improve Mandarin lip-reading performance.
@@
Visual Speech Recognition, Chinese Lip-reading, LLM
@@ I Introduction
Visual Speech Recognition (VSR), also known as lip-reading, aims to transcribe spoken language solely from silent video recordings of lip movements. This capability has broad applications in hearing-impaired communication, silent speech interfaces, and speech recognition under adverse acoustic conditions. Despite recent progress driven by deep neural networks, VSR remains a highly challenging task due to the inherent ambiguity of visual speech signals [1].
One fundamental difficulty of VSR lies in the viseme ambiguity problem: distinct phonemes may share nearly identical lip movements, while co-articulation further blurs temporal boundaries between adjacent units. As a result, purely visual observations often provide insufficient information to uniquely determine the underlying speech content. This challenge becomes significantly more severe in Mandarin Chinese, which exhibits a large number of homophones and near-homophones. As noted in prior work, lip movements in Mandarin “provide limited semantic information, particularly due to the high prevalence of visually indistinguishable syllables,” forcing models to rely heavily on contextual reasoning rather than direct visual evidence [2].
Most existing VSR systems adopt an end-to-end modeling paradigm that directly maps video frames to word or character sequences [3] [4] [5] [6]. While effective for alphabetic languages such as English, this approach is less suitable for Mandarin. Chinese is a logographic language without an explicit phoneme-level writing system, and a single syllable often corresponds to many different characters. Consequently, directly predicting characters from visual features places an excessive burden on the visual encoder to resolve phonetic and semantic ambiguities simultaneously [7].
@@
Recent studies have demonstrated that introducing explicit linguistic structure into the VSR pipeline can substantially improve recognition performance [8] [9] [10] [11]. In particular, the VALLR framework [12] proposed by Thomas et al. introduces a two-stage paradigm for English lip-reading, where a neural network first predicts phoneme sequences from video, and a large language model (LLM) subsequently decodes these phonemes into words and sentences. By explicitly modeling an intermediate phonetic representation, the LLM can leverage its extensive linguistic knowledge to recover coherent text from noisy visual predictions, achieving strong performance with reduced amounts of visual training data.
@@
However, extending this paradigm to Mandarin is non-trivial. Unlike English, Mandarin does not use a phonemic alphabet, and its romanization system, Pinyin [ikipedia:pinyin], introduces additional complexity due to tone markings and many-to-one mappings between phonetic syllables and characters. Moreover, existing LLM-based approaches typically assume correct or near-correct phoneme or Pinyin inputs. In practice, visual-only prediction of Pinyin is error-prone, and naively feeding erroneous Pinyin sequences into an LLM may amplify errors rather than correct them.
@@
Meanwhile, recent advances in automatic speech recognition (ASR) have shown that LLMs can be effectively used as post-processors for hypothesis rescoring and error correction [14] [15] [16]. By providing an LLM with N-best candidate transcriptions instead of a single hypothesis, the model can exploit linguistic context to resolve ambiguities and correct systematic recognition errors. This idea has been successfully applied in audio ASR, but its potential in the visual domain remains largely unexplored.
In this work, we propose VALLR-Pin, a Mandarin-oriented visual speech recognition framework that unifies visual modeling, phonetic representation, and language-level reasoning within a two-stage architecture. Our approach is built upon three key ideas. First, we introduce a dual-decoder VSR model that jointly predicts Mandarin characters and their corresponding Pinyin sequences, encouraging the visual encoder to learn more robust and linguistically grounded representations. Second, we construct an LLM-based refinement module that takes both the predicted Pinyin sequence and an N-best list of candidate character transcriptions as input, enabling context-aware correction of homophones and visually confusable units. Third, instead of assuming ideal phonetic inputs, we fine-tune the LLM on a dataset of model-generated errors, allowing it to adapt to the specific error patterns produced by the VSR system.
By explicitly modeling Pinyin as an intermediate representation and combining it with LLM-guided refinement, VALLR-Pin effectively bridges the gap between ambiguous visual signals and linguistically coherent Mandarin text. We expect this framework to substantially improve visual speech recognition performance, particularly in scenarios dominated by homophones and visually indistinguishable syllables.
@@ II-B2 Dual-Decoder Multi-Task Learning
where $\alpha$ is set to $0.5$ in all experiments. This multi-task formulation encourages the visual encoder to learn more robust and linguistically consistent representations.
@@ II-C2 LLM Fine-Tuning Dataset Construction
This formulation enables the LLM to explicitly learn systematic error patterns produced by the VSR model, rather than assuming ideal phonetic inputs.
@@ II-C3 LLM Fine-Tuning with LoRA
Through fine-tuning, the LLM learns to (1) map ambiguous Pinyin sequences to characters, (2) correct common visual confusion errors, and (3) exploit linguistic context to rescore and refine N-best hypotheses.
@@ III-C Experimental Setup
All VSR models share the same visual frontend consisting of 3D convolutional layers followed by a ResNet-50 backbone. The encoder is based on the SANM architecture from Paraformer. For decoding, we compare a character-only baseline with the proposed dual-decoder architecture that jointly predicts Mandarin characters and toneless Pinyin sequences.
@@
For the second-stage refinement, we fine-tune a Chinese large language model, Qwen3-4B [26], using the ms-swift framework [27]. We adopt LoRA and accelerate training with DeepSpeed [28]. The main hyperparameters are: learning_rate $1\mathrm{e}{-4}$, lora_rank =8=8, and lora_alpha =32=32. The LLM is fine-tuned on instruction-following data constructed from model-generated errors using multiple VSR checkpoints. During inference, the LLM takes the predicted Pinyin sequence and an N-best list of character hypotheses as input.
@@ III-D Comparison with Baselines
Table I compares the proposed method with baseline systems on both datasets. The CNVSRC baseline represents a strong single-decoder VSR system reported in prior work.
TABLE I: CER (%) comparison on CNVSRC-Multi.Eval and the self-collected dataset. Lower is better.
Method Setting CNVSRC-Multi.Eval Self-Collected
@@
ours (char-only) S3 30.87 38.60
@@
The proposed VALLR-Pin consistently outperforms all baselines. The improvement is more pronounced on the self-collected dataset, indicating stronger robustness under complex visual conditions.
@@ III-E Ablation Studies
Table II evaluates the impact of introducing the Pinyin decoder.
TABLE II: Ablation on dual-decoder design (CER %).
@@
Char-only Decoder 38.60
Char+Pinyin Decoder 37.23
@@
Table III compares different LLM configurations.
TABLE III: Ablation on LLM refinement (CER %).
@@
LLM (zero-shot) 37.86
LLM (fine-tuned on VSR errors) 35.49
@@
TABLE IV: Qualitative examples of transcription refinement.
@@ III-F Qualitative Analysis
Table IV presents representative examples illustrating the effect of LLM-based refinement. The LLM effectively resolves homophone errors and visually confusable characters by leveraging Pinyin guidance and contextual information.
@@ IV Conclusion
We introduced VALLR-Pin, a Pinyin-guided LLM refinement approach for Mandarin visual speech recognition. By leveraging Pinyin as an intermediate representation and adapting an LLM with error-aware instruction data, the proposed method alleviates character ambiguity in Mandarin VSR. Experimental results on the CNVSRC benchmark validate the effectiveness of the proposed framework under multi-speaker conditions.

### v2

# VALLR-Pin: Uncertainty-Factorized Visual Speech Recognition for Mandarin with Pinyin Guidance

@@ Abstract
Visual speech recognition (VSR) aims to transcribe spoken content from silent lip-motion videos and is particularly challenging in Mandarin due to severe viseme ambiguity and pervasive homophones. We propose VALLR-Pin, a two-stage Mandarin VSR framework that extends the VALLR architecture by explicitly incorporating Pinyin as an intermediate representation. In the first stage, a shared visual encoder feeds dual decoders that jointly predict Mandarin characters and their corresponding Pinyin sequences, encouraging more robust visual-linguistic representations. In the second stage, an LLM-based refinement module takes the predicted Pinyin sequence together with an N-best list of character hypotheses to resolve homophone-induced ambiguities. To further adapt the LLM to visual recognition errors, we fine-tune it on synthetic instruction data constructed from model-generated Pinyin-text pairs, enabling error-aware correction. Experiments on public Mandarin VSR benchmarks demonstrate that VALLR-Pin consistently improves transcription accuracy under multi-speaker conditions, highlighting the effectiveness of combining phonetic guidance with lightweight LLM refinement.
@@
Visual Speech Recognition, Chinese Lip-reading, LLM refinement
@@ I Introduction
Visual Speech Recognition, also known as lip-reading, aims to transcribe spoken language solely from silent video recordings of lip movements. This capability has broad applications in hearing-impaired communication, silent speech interfaces, and speech recognition under adverse acoustic conditions. Despite recent progress driven by deep neural networks [1] [2], VSR remains a highly challenging task due to the inherent ambiguity of visual speech signals.
One fundamental difficulty of VSR lies in the viseme ambiguity problem: where distinct phonemes may share nearly identical lip movements, while co-articulation further blurs temporal boundaries between adjacent units. In Mandarin, this challenge is further exacerbated by a mismatch between visual discriminability and linguistic representation, rendering Mandarin VSR inherently ill-posed. While visual cues can often constrain speech content at the syllable level, directly predicting characters from visual input requires resolving a one-to-many mapping, in which visually indistinguishable syllables may correspond to a large set of characters [3, 4]. This intrinsic ambiguity suggests that explicitly modeling an intermediate syllable-level representation is a more natural way to factorize uncertainty in Mandarin VSR.
@@
Most existing VSR systems adopt an end-to-end modeling paradigm that directly maps video frames to word or character sequences [5] [6] [7] [8]. Recent studies have demonstrated that introducing explicit linguistic structure into the VSR pipeline can substantially improve recognition performance. [9] [10] [11] [12]. In particular, the VALLR framework [13] proposed by Thomas et al. introduces a two-stage method for English lip-reading, where a neural network first predicts phoneme sequences from video, and a large language model subsequently decodes these phonemes into sentences. By explicitly modeling an intermediate phonetic representation, the framework allows the LLM to leverage linguistic priors to recover coherent text from noisy visual predictions, achieving strong performance with reduced amounts of visual training data.
@@
However, unlike English, Mandarin does not use a phonemic alphabet, and its romanization system, Pinyin [14], introduces additional complexity due to tone markings and many-to-one mappings between phonetic syllables and characters. Moreover, most existing LLM-based approaches implicitly assume relatively accurate phoneme or Pinyin inputs . In practice, visual-only prediction of Pinyin remains error-prone, and naively feeding erroneous phonetic sequences into an LLM may amplify recognition errors rather than correct them.
@@
Meanwhile, recent advances in automatic speech recognition (ASR) have shown that LLMs can be effectively used as post-processors for hypothesis rescoring and error correction [15] [16] [17]. By providing an LLM with N-best candidate transcriptions instead of a single hypothesis, the model can exploit linguistic context to resolve ambiguities and correct systematic recognition errors. Although this paradigm has shown promise in audio-based ASR, its application to visual speech recognition remains largely underexplored.
In this work, we propose VALLR-Pin, a Mandarin-oriented visual speech recognition framework that unifies visual modeling, phonetic representation, and language-level reasoning within a two-stage architecture. Instead of directly predicting characters from visual features, we reformulate Mandarin VSR as an uncertainty-factorized inference problem by jointly modeling syllable-level and character-level decoding processes, thereby explicitly decomposing visual uncertainty and linguistic ambiguity. On top of this factorized representation, we incorporate an LLM-based refinement module that operates under phonetic and candidate-level constraints to correct residual errors, rather than acting as an independent recognizer. Furthermore, to mitigate error propagation from imperfect phonetic predictions, we fine-tune the LLM using model-generated error patterns, enabling robust adaptation to the characteristic uncertainties of visual speech recognition.
@@ II-B2 Dual-Decoder Multi-Task Learning
where $\alpha$ is set to $0.5$ in all experiments. This multi-task formulation encourages the visual encoder to learn more robust and linguistically grounded representations.
@@ II-C2 LLM Fine-Tuning Dataset Construction
This formulation allows the LLM to better capture systematic error patterns produced by the VSR model.
@@ II-C3 LLM Fine-Tuning with LoRA
Through fine-tuning, the LLM is guided to (1) map ambiguous Pinyin sequences to characters, (2) correct common visual confusion errors, and (3) exploit linguistic context to rescore and refine N-best hypotheses.
@@ III-A Datasets
CMLR
This is a Chinese sentence-level lip-reading dataset, consisting of 11 speakers and 102,072 spoken sentences [23]. The model proposed by Pingchuan Ma et al. achieves a character error rate of 9.10 on this dataset [24].
@@ III-C Experimental Setup
We build a char-only baseline model by integrating the Paraformer structure into the CNVSRC 2025 baseline model. All VSR models share the same visual frontend consisting of 3D convolutional layers followed by a ResNet-50 backbone. The encoder is based on a SANM Encoder. For decoding, the character-only baseline employs a SANM decoder; on this basis, our improved model introduces an additional Transformer decoder-based Pinyin decoder, enabling it to simultaneously predict Mandarin characters and toneless Pinyin sequences.
@@
For the second-stage refinement, we fine-tune a large language model, Qwen3-4B-Instruct-2507 [29]  [^1] https://modelscope.cn/models/Qwen/Qwen3-4B-Instruct-2507, using the ms-swift framework [30]. We adopt LoRA and accelerate training with DeepSpeed [31]. The main hyperparameters are: learning_rate $1\mathrm{e}{-4}$, lora_rank =8=8, and lora_alpha =32=32. The LLM is fine-tuned on instruction-following data constructed from model-generated errors using multiple VSR checkpoints. During inference, the LLM takes the predicted Pinyin sequence and an N-best list of character hypotheses as input.
@@ III-D Comparison with Baselines
Table I compares the proposed method with baseline systems on both datasets.
TABLE I: CER (%) comparison on CNVSRC-Multi.Dev and the self-collected dataset. Lower is better.
Method Setting CNVSRC-Multi.Dev Self-Collected
@@
Char-only Baseline S3 30.87 38.60
@@
TABLE II: CER (%) comparison on CMLR. Lower is better.
Model CER
Ma’s model 9.10
+ Pinyin-Guided Refinement 7.89
The char-only baseline achieves a lower CER than the CNVSRC 2025 competition baseline. Additionally, the proposed VALLR-Pin consistently outperforms all baselines. The improvement is more pronounced on the self-collected dataset, indicating stronger robustness under complex visual conditions.
Additionally, we conduct a comparative experiment with Pingchuan Ma’s model on the CMLR dataset; integrating the dual-decoding structure and two-stage LLM Refinement into their model leads to a further reduction in CER, with results presented in Table II.
@@ III-E Ablation Studies
Table III evaluates the impact of introducing the Pinyin decoder.
TABLE III: Ablation on dual-decoder design (CER %).
@@
Char-only Baseline 38.60
+ Pinyin Decoder 37.23
@@
Table IV compares different LLM configurations.
TABLE IV: Ablation on LLM refinement (CER %).
@@
+ Zero-shot LLM 37.86
+ LLM Refinement 35.49
@@
TABLE V: Qualitative examples of transcription refinement.
@@ III-F Qualitative Analysis
Table V presents representative examples illustrating the effect of LLM-based refinement. The LLM effectively resolves homophone errors and visually confusable characters by leveraging Pinyin guidance and contextual information.
@@ IV Conclusion
We presented VALLR-Pin, an uncertainty-aware framework for Mandarin visual speech recognition that explicitly factorizes visual ambiguity through an intermediate syllable-level representation. By decomposing Mandarin VSR into coupled syllable-level and character-level inference processes, the proposed approach addresses the inherently ill-posed nature of direct character prediction from visual input. On top of this factorized representation, a lightweight LLM-based refinement module is employed as a complementary constraint mechanism to correct residual linguistic ambiguities, rather than acting as an independent recognizer. Experimental results on the CNVSRC benchmark demonstrate consistent improvements under multi-speaker conditions, indicating that uncertainty-aware modeling combined with constrained language-level refinement provides an effective and modular enhancement to visual speech recognition systems.